At General Electric, shift 1 produced 4 times as much as shift 2. General Electric’s total production for the month of July was 5,500 jet engines. What was the output for each shift?
Answer by checkley77(12569)

4X+X=5,500
5X=5,500
X=5,500/5
X=1,100 ANS. FOR THE OUTPUT FOR THE SHIFT 2.
5,500-1,100=4,400 ANS. FOR THE OUTPUT FOR THE SHIFT 1.